Mal87:

A horse begins 10 meters along a road from an intersection. After 6 seconds it is 80 meters further west. What is the average velocity of the horse? A. about 13.3 m/s east B. about 13.3 m/s west C. about 13.8 m/s east D. about 13.8 m/s west

Mal87:

@Shadow I know you have to divide but I don't know what to divide.

Shadow:

So it is asking us for velocity, which depends on both speed and direction (since it's a vector property). Let's start off with what direction is the horse going in?

Mal87:

West I think

Shadow:

Correct. If it is 'further west' than it was prior to the six seconds, then it is definitely traveling in the westward direction. Next, what is the velocity? The formula for velocity is the following? \[\Delta V = \frac{ \Delta x }{ \Delta t }\] Have you seen this formula before?

Mal87:

No I haven't

Shadow:

Let me rewrite it real quick. \[V _{avg} = \frac{ \Delta x }{ \Delta t}\] The triangle sign (Delta) stands for change in. Change in x is something we call displacement. t stands for time, so delta t stands for change in time. Formula for displacement: \[\Delta x = x _{f} - x_{i}\] Where you take the final x coordinate and subtract it by the initial x coordinate. So for this problem, in order to solve for the displacement. We know that he has traveled 80 meters further west, so that is our displacement. As for change in time, that is six seconds. \[V _{avg} = \frac{ 80 }{ 6 } = 13.3m/s \] In the westward direction

Mal87:

Thank you for helping me I will copy the formula so I have it if I need it again and use this problem as an example.

Shadow:

Okay cool, and no problem